# itemizestditemize

### v1

@@ III Resource-Aware Distributed Coordination
We define the problem of Resource-Aware Distributed Coordination (1) with regard to the action coordination module in Fig. 2. To this end, we use the notation:
@@
We also lay down the following framework for the robot-to-robot communication network and the objective function that captures the multi-robot task at hand.
Communication Neighborhood. At the beginning of each planning step (Fig. 2), given the observed environment and state of the robots, the robots decide with which others to establish communication, subject to their onboard bandwidth constraints. Specifically, we assume that each robot $i$ can receive information from up to $\alpha_{i}$ other robots due to onboard bandwidth constraints. Thus, it must be $|{\cal N}_{i}|\;\leq\alpha_{i}$. In Section V, we provide theoretical characterizations that inform how the robots may select their neighbors to optimize the coordination performance. In our simulations over a scenario of road coverage, the robots select their neighbors based on physical proximity, as justified by the results in Section V.
@@
Communication Data Rate. All communication channels $(j\rightarrow i)\in{\cal E}$ have finite data rate, i.e., communication speed. In the simulations, we assume the data rate is $10$MBytes/sec, in accordance with typical v2v communication speeds [38]. Due to the finite data rate, the decision time of action coordination depends on both (i) the number of communication rounds and (ii) the size of transmitted messages it requires for the robots to find a joint plan. Without loss of generality, we assume all communication channels have the same data rate in this paper.
Objective Function. The robots coordinate their actions to maximize an objective function. In active information gathering tasks, such as area coverage, target tracking, and persistent monitoring, typical objective functions are the covering functions [39, 20, 40]. These functions capture how much area/information is observed given the actions of all robots. They satisfy the properties below (Definitions 1 and 2).
@@
$f(s)-f(s\,|\,{\cal A})\geq f(s\,|\,{\cal B})-f(s\,|\,{\cal A},\,{\cal B}),$ (2)
for any disjoint ${\cal A}\subseteq{\cal V}$ and ${\cal B}\subseteq{\cal V}$ (${\cal A}\cap{\cal B}=\emptyset$) and $s\in{\cal V}$.
The 2nd-order submodularity is another intrinsic property to the function. Intuitively, if $f({\cal A})$ captures the area covered by a set ${\cal A}$ of cameras, then marginal gain of the marginal gains drops when more cameras are already activated.
Problem 1 (Resource-Aware Distributed Coordination).
Each robot $i\in{\cal N}$ independently selects an action $a_{i}$, using only information from and about its neighbors ${\cal N}_{i}$, such that the robot actions $\{a_{i}\}_{i\,\in\,{\cal N}}$ jointly solve the optimization problem [^1] In the Appendix, we extend our framework beyond 2nd-order submodular functions, and present results for functions that are merely submodular or approximately submodular.
@@
1 is resource-aware in that it requires each robot to coordinate actions with its neighbors only, receiving information only about them instead of more robots in the network. The reason is to curtail the explosion of information passing across the network and, thus, to enable rapid decision-making. This is in contrast to standard distributed methods that allow information about the whole network to travel to all other robots via information passing, i.e., multi-hop communication [21, 39, 43]. Multi-hop communication does not reduce the amount of information flowing in the network compared to centralized methods, introducing impractical communication delays, as we will demonstrate in the simulations.
In the resource-aware framework of 1, the network configuration affects the achievable maximal value of $f$ in 3 since the network configuration decides what information becomes available to each of the robots. For example, in a fully decentralized network where all robots select actions in isolation, the achievable value of $f$ will be lower than the achievable value in a fully centralized network where all robots coordinate with each other and jointly select actions. Similarly, consider a scenario where the robots are identical mobile cameras that start from the same location and have the same options to move around, and where $f$ is the union of the area covered by the robots’ FOVs. Then, in the fully decentralized setting, all robots may end up looking at the same area, due to the lack of any coordination. Thus, $f$ will be equal to the area covered by one robot since all robots’ views overlap. In contrast, in the fully centralized setting, the robots can coordinate to avoid overlapping their FOVs, thus maximizing the total covered area $f$.
In this paper, we provide an analysis of how the network topology at the local level —each robot’s neighborhood ${\cal N}_{i}$— affects the near-optimality of coordination at the global level. The analysis applies over the spectrum from fully decentralized to fully centralized networks, offering insights into how the robots may tune the sparsity of the network to balance the trade-off between decision speed and near-optimality.
Along with the theoretical characterizations above, we also provide a communication-efficient distributed algorithm that requires only a few coordination rounds and exchanging short messages among the robots, per the limited information-sharing setting of 1, and that is up to two to three orders faster than the state-of-the-art near-optimal algorithms. The algorithm is presented next.
@@ IV Resource-Aware distributed Greedy (RAG) Algorithm
Figure 4: RAG vs. Sequential Greedy (SG[18]. The two algorithms are compared in their execution steps and in the time they need to terminate when they are run by five agents in two scenarios with the communication networks being: (i) an undirected line graph and (ii) an undirected star graph, respectively. In the figure, nodes $1$ to $5$ are the agents $1$ to $5$; black lines are undirected communication links; $\{a_{i},g_{i}\}$ denotes agent $i$’s newly updated action candidate $a_{i}$ with marginal gain $g_{i}$; $\{a_{i}\}$ and $\{g_{i}\}$ alongside red arrows are the actions and marginal gain values being transmitted between two agents; the transparent nodes are the agents that have selected their actions and thus already ended running RAG; and the transparent edges are the communication channels that have become “disappeared” since at least one end of them has finished RAG. The implementation of RAG ((a)–(d) and (f)–(i)) results in way shorter decision time than SG ((e) and (j)) in both scenarios with the line and star graphs, respectively.
@@ IV-B Comparison to the Sequential Greedy algorithm (SG)
RAG is compared with SG [18] in Fig. 4. We rigorously present SG next, and provide a qualitative comparison with RAG. Rigorous comparisons in terms of runtime and approximation performance are postponed to Sections V and VI.
@@
The above action-selection features of RAG —parallelization and action selection before all neighbors have chosen an action— can speed up the algorithm’s termination, as we rigorously present in Section VI. Also, they enable RAG to work on arbitrary communication topologies, a fact that can further contribute to faster runtimes. For example, SG requires a line path connecting all robots in the action-selection sequence. If such a path does not exist (see star graph example in Fig. 4), the $i$-th robot in the action-selection sequence cannot necessarily communicate directly with the $(i+1)$-th robot. Then, SG requires extra communication rounds for message relaying, further delaying its termination. Specifically, given the limited communication speed of robot-to-robot communication channels [38], the termination of SG is delayed due to both the increased number of communication rounds, and the communication delay incurred from relaying the actions of multiple robots —the robots in the sequence that have chosen an action so far— across the network.
@@ V Approximation Guarantees:
We present the a priori and posteriori suboptimality bounds of RAG (Theorems 1 and 2). They both capture the suboptimality cost due to decentralization, that is, due to each agent coordinating only with a few agents —its neighbors— and receiving information only about them, in favor of decision speed.
The two bounds are useful as follows: The a priori bound enables each agent to design its neighborhood to minimize its local suboptimality cost subject to its communication-bandwidth constraints. The a posteriori bound results in two practical observations: first, decentralized coordination can marginally be as near-optimal as centralized coordination (Remark 7); second, larger neighborhoods do not necessarily result in better coordination performance when we account for both the decision time and the suboptimality of the algorithm (Remark 8). We validate and leverage the above in the experimental results of Section VII.
In the following paragraphs, we first introduce the novel notion of Centralization of Information to quantify the a priori suboptimality bound (Section V-A). Then, we present first the a priori bound of RAG that involves the proposed novel notion (Section V-B) and second the a posteriori bound (Section V-C). Finally, we compare the a priori bound with the a priori bounds in the state of the art (Section V-D).
@@ V-A Centralization of Information
We introduce the notion of centralization of information (${\sf coin}$). In Section V-B, we use the notion to quantify the suboptimality cost due to decentralization. Particularly, ${\sf coin}$ measures how the agents’ actions overlap due to the agents not coordinating with their non-neighbors. We also relate ${\sf coin}$ to the classical notion of total curvature [44] and to pair-wise consistency [39] (Remarks 2 and 3), and show that ${\sf coin}$ is a less conservative measure of action overlap.
@@
  • The total curvature [44] of a function $f\colon 2^{{\cal V}_{\cal N}}\mapsto\mathbb{R}$ that is non-decreasing and submodular, and that, without loss of generality, $f(a)\neq 0$, for any agent’s action $a$, i.e., for any $a\in\bigcup_{i\,\in\,{\cal N}}{\cal V}_{i}$, is defined as:
@@
    $\kappa_{f}$ measures how the action $a$ of an agent can overlap with the actions of all other agents in the worst case. Particularly, $\kappa_{f}\in[0,1]$, and if $\kappa_{f}=0$, then $f({\cal A})-f({\cal A}\setminus\{a\})=f(a)$, for all $a\in{\cal A}$, i.e., the action $a$ of an agent does not overlap with the actions of any other agents. In contrast, if $\kappa_{f}=1$, then there exists $a\in{\cal A}$ such that $f({\cal A})=f({\cal A}\setminus\{a\})$, i.e., the agent with the action $a$ has no contribution to $f({\cal A})$ in the presence of all other agents.
@@
${\sf coin}_{f,i}$ measures how much the action $a_{i}$ of the agent $i$ can be substituted from the actions $\{a_{j}\}_{j\,\in\,{\cal N}_{i}^{c}}$ of its non-neighbors. In the best case where $a_{i}$ cannot be substituted at all, i.e., $f(a_{i}\,|\,\{a_{j}\}_{j\,\in\,{\cal N}_{i}^{c}})=f(a_{i})$, then indeed ${\sf coin}_{f,i}=0$. In the worst case instead where $a_{i}$ is fully substituted, i.e., $f(a_{i}\,|\,\{a_{j}\}_{j\,\in\,{\cal N}_{i}^{c}})=0$, then indeed ${\sf coin}_{f,i}=f(a_{i})$.
From an information-theoretic perspective, ${\sf coin}_{f,i}$ measures how much the information collected by $a_{i}$ overlaps with the information collected by $\{a_{j}\}_{j\,\in\,{\cal N}_{i}^{c}}$. Rigorously, if $f$ is an entropy metric, then ${\sf coin}_{f,i}$ is mutual information [45]. Thus, in this context, ${\sf coin}_{f,i}=0$ if and only if the information collected by $a_{i}$ is decentralized from (independent of) the information collected by $\{a_{j}\}_{j\,\in\,{\cal N}_{i}^{c}}$. In this sense, ${\sf coin}_{f}$ captures the decentralization of information across the network.
Remark 2 (Relation to Total Curvature [44]).
@@
The sum of all $\{{\sf coin}_{f,i}\}_{i\,\in\,{\cal N}}$ will be used in the next section to characterize the global suboptimality cost due to decentralization. Given this characterization, we may want to enable the agents to pick their neighborhoods to minimize coin subject to their communication-bandwidth constraints. But observe that ${\sf coin}_{f}$ can be uncomputable a priori since, in distributed settings, agent $i$ will not have access a priori to the actions of its non-neighbors, and may not even know who are its non-neighbors since it may not know how many agents exist in the network beyond its neighbors. Notwithstanding, finding a computable upper bound for ${\sf coin}_{f,i}$ may be easy, as we demonstrate in the following example.
@@
The upper bound in eq. 8 as a function of the communication range $r_{i}$ is plotted in Fig. 5(c). As expected, it tends to zero for increasing $r_{i}$. Particularly, when the distance of agent $i$ for its nearest non-neighbor is larger than $2r_{s}$, then the FOVs of agent $i$ and its non-neighbors cannot overlap, thus ${\sf coin}_{f,i}=0$.
@@
RAG (this paper) ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\leq\tau_{f}\,\max_{i\,\in\,{\cal N}}{(|\mathcal{V}_{i}|\,|\mathcal{N}_{i}|)}}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\leq(\tau_{c}+\tau_{\#})\,(|{\cal N}|-1)}$ ${1}/{2}\left[{\sf OPT}-{\sum_{i\in{\cal N}}{{\sf coin}_{f,i}({\cal N}_{i})}}\right]$, $\xi\,{\sf OPT},\;\xi\in\left[\frac{1-\kappa_{f}}{2-\kappa_{f}},\frac{1}{2}\right]$ possibly disconnected, directed

### v2

@@ III Resource-Aware Distributed Submodular Optimization
We define the problem of Resource-Aware Distributed Submodular Optimization (1) with regard to the action coordination module in Fig. 2. To this end, we use the notation:
  • ${\cal N}$ is the set of robots;
@@
We also use the following framework. For easiness of illustration, we present the framework focusing on multi-robot information-gathering tasks such as simultaneous localization and mapping (SLAM), target tracking, surveillance [1, 15, 17]. Nonetheless, 1 applies to any distributed optimization problem with a submodular objective function.
Robot Dynamics. We assume mobile robots that act as mobile sensors. Their motion dynamics may take the form:
$x_{i,t}=f_{i}(x_{i,t-1},u_{i,t-1})+n_{i,t},\quad t=1,2,\ldots,$ (1)
where $x_{i,t}\in\mathbb{R}^{n_{x_{i,t}}}$ denotes the state of robot $i$ at time step $t$, $u_{i,t}\in\mathcal{U}_{i,t}$ denotes the control input applied to robot $i$, where $\mathcal{U}_{i,t}$ is a finite set of admissible control inputs [1], and $n_{i,t}$ is process noise, e.g., Gaussian noise.
Target Dynamics. The robots use their sensors to measure and infer information about the environment. For example, in SLAM and target tracking, at each time step $t$, the robots aim to estimate a target state vector $y_{t}$ that encapsulates the landmark locations or the mobile targets’ to be localized [1]. Generally, $y_{t}$ evolves with motion dynamics of the form:
$y_{t}=\phi(y_{t-1})+w_{t},\quad t=1,2,\ldots,$ (2)
where $w_{t}$ denotes process noise, e.g., Gaussian noise.
Sensor Model. For each robot $i$, we assume sensor model:
$z_{i,t}=g_{i}(x_{i,t},y_{t})+v_{i,t}(x_{i,t}),\quad t=1,2,\ldots,$ (3)
where $z_{i,t}$ denotes the measurement taken by robot $i$ at time $t$, and $v_{i,t}$ is noise that possibly depends on the robots’ and targets’ state. For example, $v_{i,t}$ may be Gaussian noise with mean $\mu_{v_{i,t}}(x_{i,t},y_{t})$, and covariance $\Sigma_{v_{i,t}}(x_{i,t},y_{t})$.
Communication Neighborhood. Before each action coordination step (Fig. 2), given the observed environment and states of the robots, the robots decide with which others to establish communication, subject to their onboard bandwidth constraints. Specifically, we assume that each robot $i$ can receive information from up to $\alpha_{i}$ other robots due to onboard bandwidth constraints ($|{\cal N}_{i}|\;\leq\alpha_{i}$).
@@
Communication Data Rate. All communication channels $(j\rightarrow i)\in{\cal E}$ have finite data rates (communication speeds). In the simulations, we assume two values for the data rates: $0.25$ Mbps (simulating the Digi XBee 3 Zigbee 3.0 antenna), and $100.0$ Mbps (simulating the Silvus SL5200 antenna). Due to the finite data rates, the decision time of action coordination depends on both (i) the number of communication rounds and (ii) the size of transmitted messages it requires for the robots to find a joint plan. We assume that all communication channels have the same data rate in this paper.
Objective Function. At each action coordination step, the robots choose actions to maximize an objective function $f$ that captures the current multi-robot task. For example, in SLAM and target tracking, at each time step $t$, $f$ can be [1]:
$f(\{a_{i}\}_{i\,\in\,{\cal N}})=\sum_{\tau=t+1}^{t+T}\mathbf{h}(y_{t}|\{z_{i,\;t+1:t+T}(a_{i})\}_{i\,\in\,{\cal N}}),$ (4)
per the robot, sensing, and target models in eqs. 1, 3 and 2,
$a_{i}\triangleq\{u_{i,t+1},\dots,u_{i,t+T}\},$
$z_{i,t+1:t+T}\triangleq\{z_{i,t+1},\ldots,z_{i,t+T}\}$, $T$ is an action-planning look-ahead horizon, and $\mathbf{h}(\cdot|\cdot)$ is an information metric such as the conditional entropy [1] or the mutual information [15]. When $\mathbf{h}$ is conditional entropy and the process and sensor noises in eqs. 1, 3 and 2 are uncorrelated Gaussian, then eq. 4 is computable a priori given any action set $\{a_{i}\}_{i\,\in\,{\cal N}}$: $\mathbf{h}$ is then equal to the $\log\det$ volume of the Kalman filtering uncertainty over the horizon $T$ and is, thus, independent of the measurements’ realization [1].
Condition entropy, mutual information, and, more broadly, covering functions [1, 15, 44, 30, 45] are used to model active information-gathering tasks such as target tracking, SLAM, and surveillance [1, 15, 17]. These functions capture how much information is observed given the actions of all robots. They belong to a broader class of functions that satisfy the properties below (Definitions 1 and 2).
@@
$f(s\,|\,{\cal C})-f(s\,|\,{\cal A}\cup{\cal C})\geq f(s\,|\,{\cal B}\cup{\cal C})-f(s\,|\,{\cal A}\cup{\cal B}\cup{\cal C}),$ (5)
for any disjoint ${\cal A},{\cal B},{\cal C}\subseteq{\cal V}$ (${\cal A}\cap{\cal B}\cap{\cal C}=\emptyset$) and $s\in{\cal V}$.
The 2nd-order submodularity is another intrinsic property of the function. Intuitively, if $f({\cal A})$ captures the area covered by a set ${\cal A}$ of cameras, then the marginal gain of marginal gains drops when more cameras are already activated.
Problem 1 (Resource-Aware Distributed Submodular Optimization).
Consider a normalized and non-decreasing submodular function $f\colon 2^{\prod_{i\in{\cal N}}{\cal V}_{i}}\mapsto\mathbb{R}$ that captures the optimization task at hand at the current action coordination step. Each robot $i\in{\cal N}$ selects an action $a_{i}\in{\cal V}_{i}$, using only information from and about its neighbors ${\cal N}_{i}$, such that the actions $\{a_{i}\}_{i\,\in\,{\cal N}}$ jointly solve the optimization problem: [^1] We extend our framework to any normalized, non-decreasing and merely submodular or approximately submodular function $f$ in the Appendix.
@@
Remark 1 (Generality).
1 applies to any distributed optimization problem in control, robotics, machine learning, and beyond where the objective function is non-decreasing and submodular, including the information-gathering tasks captured via eqs. (1) to (4). [^2] We refer the reader to the papers [1, 15] for implementation details on rapidly computing single-robot actions $a_{i}\triangleq\{u_{i,t+1},\dots,u_{i,t+T}\}$ that optimize $f$ per eqs. (1) to (4), given actions for the remaining robots.
1 is resource-aware in that it requires each robot to coordinate actions with its neighbors only, receiving information only about them instead of more robots in the network. The reason is to curtail the explosion of information passing across the network and, thus, to enable rapid coordination. This is in contrast to standard distributed methods that allow information about the whole network to travel to all other robots via information passing (multi-hop communication) [31, 44, 48]. Multi-hop communication does not reduce the amount of information flowing in the network compared to centralized methods, introducing impractical communication delays, as we will demonstrate in the simulations.
@@ IV Resource-Aware distributed Greedy (RAG) Algorithm
Figure 4: RAG vs. Sequential Greedy (SG[11]. The two algorithms are compared in their execution steps and in the time they need to terminate. We present the case of five agents in two scenarios with the communication networks being (i) an undirected line graph (top row), and (ii) an undirected star graph (bottom row). Nodes $1$ to $5$ are the agents $1$ to $5$; black lines are undirected communication links; $\{a_{i},g_{i}\}$ denotes agent $i$’s newly updated action candidate $a_{i}$ with marginal gain $g_{i}$; $\{a_{i}\}$ and $\{g_{i}\}$ alongside red arrows are the actions and marginal gain values being transmitted between two agents; the transparent nodes are the agents that have selected their actions and thus already ended running RAG; and the transparent edges are the communication channels that have become “disappeared” since at least one end of them has finished RAG. The implementation of RAG ((a)–(d) and (f)–(i)) results in shorter decision time than SG ((e) and (j)) in both scenarios with the line and star graphs, respectively.
@@ IV-B Comparison to the Sequential Greedy algorithm (SG)
RAG is compared with SG [11] in Fig. 4. We rigorously present SG next, and provide a qualitative comparison with RAG. The rigorous comparison of runtime and approximation performance is postponed to Sections V and VI, where, e.g., we prove that RAG’s runtime scales linearly with the number of the robots whereas SG’s scales cubically.
@@
The above action-selection features of RAG —parallelization and action selection before all neighbors have chosen an action— can further speed up the algorithm’s termination. Also, they enable RAG to work on arbitrary communication topologies. In contrast, SG requires a line path connecting all robots in the action-selection sequence. If such a path does not exist (see the star graph example in Fig. 4), the $i$-th robot in the action-selection sequence cannot communicate directly with the $(i+1)$-th robot. Then, SG requires extra communication rounds for message relaying, further delaying its termination. Specifically, given the limited communication speed of robot-to-robot communication channels [49], the termination of SG is delayed due to both the increased number of communication rounds, and the communication delay incurred from relaying the actions of multiple robots —the robots in the sequence that have chosen an action so far— across the network.
@@ V Approximation Guarantees:
We present a priori and posteriori suboptimality bounds for RAG (Theorems 1 and 2). To this end, we first introduce the notion of centralization of information to quantify the a priori bound (Section V-A). Then, we present the a priori bound (Section V-B) and the a posteriori bound (Section V-C). Finally, we compare the a priori bound with the a priori bounds in the state of the art (Section V-D).
@@ V-A Centralization of Information
We introduce the notion of centralization of information (${\sf coin}$). We use the notion to quantify the a priori suboptimality cost due to decentralization. ${\sf coin}$ measures how the agents’ actions can overlap due to not coordinating with their non-neighbors. We also relate ${\sf coin}$ to curvature [50] and pair-wise consistency [44] (Remarks 3 and 4), and show that ${\sf coin}$ is a less conservative measure of action overlap.
@@
Definition 3 (Curvature [50]).
Consider a function $f:2^{{\cal V}_{\cal N}}\to\mathbb{R}$ that is non-decreasing and submodular. Without loss of generality, we assume that for any action $a\in\{\mathcal{V}_{i}\}_{i\in{\cal N}}$, it holds that $f({a})\neq 0$. The curvature of $f$ is defined as:
@@
$\kappa_{f}$ measures how much an agent’s action $a$ can overlap with other agents’ actions. Particularly, $\kappa_{f}\in[0,1]$, and if $\kappa_{f}=0$, then $f({\cal A})-f({\cal A}\setminus\{a\})=f(a)$, for all $a\in{\cal A}$, i.e., no agent’s action overlaps with any other agent’s actions. In contrast, if $\kappa_{f}=1$, then there exists an action $a\in{\cal A}$ such that $f({\cal A})=f({\cal A}\setminus\{a\})$, i.e., action $a$ has no contribution to $f({\cal A})$ in the presence of all other agents.
@@
${\sf coin}_{f,i}$ measures how much $a_{i}$ can overlap with the actions of agent $i$’s non-neighbors. In the best case where $a_{i}$ does not overlap at all, i.e., $f(a_{i}\,|\,\{a_{j}\}_{j\,\in\,{\cal N}_{i}^{c}})=f(a_{i})$, then ${\sf coin}_{f,i}=0$. In the worst case instead where $a_{i}$ is fully overlapped, i.e., $f(a_{i}\,|\,\{a_{j}\}_{j\,\in\,{\cal N}_{i}^{c}})=0$, then ${\sf coin}_{f,i}=f(a_{i})$.
From an information-theoretic perspective, ${\sf coin}_{f,i}$ measures how much the information collected by $a_{i}$ overlaps with the information collected by $\{a_{j}\}_{j\,\in\,{\cal N}_{i}^{c}}$. Rigorously, if $f$ is an entropy metric, then ${\sf coin}_{f,i}$ is mutual information [51]. Thus, ${\sf coin}_{f,i}=0$ if and only if the information collected by $a_{i}$ is decentralized from (independent of) the information collected by $\{a_{j}\}_{j\,\in\,{\cal N}_{i}^{c}}$. In this sense, ${\sf coin}_{f}$ captures the decentralization of information across the network.
Remark 3 (Relation to Curvature [50]).
@@
The sum of all $\{{\sf coin}_{f,i}\}_{i\,\in\,{\cal N}}$ will be used in the next section to characterize the global suboptimality cost due to decentralization. Given this characterization, we may want to enable the agents to pick their neighborhoods to minimize coin subject to their communication-bandwidth constraints. ${\sf coin}_{f}$ can be uncomputable a priori since agent $i$ will not have access to the actions of its non-neighbors. Notwithstanding, finding a computable upper bound for ${\sf coin}_{f,i}$ may be easy, as we demonstrate in the following example.
@@
The bound as a function of the communication range $r_{i}$ is plotted in Fig. 5(c). It tends to zero for increasing $r_{i}$, as expected. When the distance of agent $i$ for its nearest non-neighbor is larger than $2r_{s}$, then the FOVs of agent $i$ and its non-neighbors cannot overlap, thus ${\sf coin}_{f,i}=0$.
@@
RAG (this paper) ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\leq\tau_{f}\,\max_{i\,\in\,{\cal N}}{(|\mathcal{V}_{i}|\,|\mathcal{N}_{i}|)}}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\leq(\tau_{c}+\tau_{\#})\,(|{\cal N}|-1)}$ ${1}/{2}\left[{\sf OPT}-{\sum_{i\in{\cal N}}{{\sf coin}_{f,i}({\cal N}_{i})}}\right]$, $\xi\,{\sf OPT},\;\xi\in\left[1-\kappa_{f},\frac{1}{1+\kappa_{f}}\right]$ possibly disconnected, directed